I have to create and solve an equation to answer this question. The perimeter of a rectangle is 210m. The length is 7m longer than the width. Determine the dimensions of the rectangle.
Answer by checkley77(12844)

L=W+7
P=2L+2W
210=2(W+7)+2W
210=2W+14+2W
210=4W+14
4W=210-14
4W=196
W=196/4
W=49 ANS. FOR THE WIDTH.
L=49+7=56 ANS. FOR THE LENGTH.
PROOF:
210=2*56+2*49
210=112+98
210=210